# Generative models for crystalline materials

### v1

@@ 6 Summary and outlook
Beyond models and representations, available data is a key ingredient for further progress. Most early generative models were trained on simpler subsets of the Materials Project, e.g., perov or MP20. Recently, benchmarks also include more challenging datasets with larger unit cells, e.g., MPTS-52. As only 10% of the structures in the Materials Project have more than 52 atoms, high metrics on benchmarks like MPTS-52 will cover most of the application-relevant materials regarding unit cell size, but still exclude specific materials classes with larger unit cells, such as metal-organic frameworks. Scaling to larger unit cell sizes benefits from models that work on the asymmetric unit representation. It is currently unclear whether transformer-based architectures with specific tokenization related to asymmetric unit cells will have an advantage over GNN-based models, which typically require full connectivity among all atoms in the asymmetric unit cell.

### v2

@@ 6 Summary and outlook
Beyond models and representations, available data is a key ingredient for further progress. Most early generative models were trained on simpler subsets of the Materials Project, e.g., perov or MP-20. Recently, benchmarks also include more challenging datasets with larger unit cells, e.g., MPTS-52. As only 10% of the structures in the Materials Project have more than 52 atoms, high metrics on benchmarks like MPTS-52 will cover most of the application-relevant materials regarding unit cell size, but still exclude specific materials classes with larger unit cells, such as metal-organic frameworks. Scaling to larger unit cell sizes benefits from models that work on the asymmetric unit representation. It is currently unclear whether transformer-based architectures with specific tokenization related to asymmetric unit cells will have an advantage over GNN-based models, which typically require full connectivity among all atoms in the asymmetric unit cell.